what is the answer for this equation 1/3x-1/5x=4/3
Answer by pirategirl(1)

1/3x - 1/5x= 4/3
5/15x - 3/15x= 4/3
45/60x = 4/3
3/4x = 4/3
4/3 * 3/4x = 4/3
12/12x = 4/3
1 x = 4/3
x = 4/3
you have to find the common denominator first.
1/3x converts to 5/15x
1/5x converts to 3/15x
then cross multiply 3/15x times 5/15x
3 times 15 = 45
5 times 15 = 60
this reduces to 3/4x
to separate 3/4 from x you take the inverse of 3/4 and multiply
now you have 3/4 times 4/3 x which is 12/12x which is the same as 1 times x
whatever you do to one side you must do to the other so 4/3 times 1 is equal to 4/3
I hope this helps!